If we have a binomial probability distribution in which n= 23 and p= .64, what is its standard deviation?
Answer #1

solution:

p = 0.64

q = 1 - p =1-0.64=0.36

n = 23

Using binomial distribution,

Mean = = n * p = 23*0.64=14.72

Standard deviation = =  n * p * q =  23*0.64*0.36=2.30